OpenStudy (anonymous):

cot theta = -1/10. Where theta lies in Quadrant 2. What is csc theta

OpenStudy (anonymous):

cot = x/y csc = r/y =sqrt(101)/10

OpenStudy (anonymous):

r = sqrt(1^2+10^2) = sqrt 101 For proof.

OpenStudy (anonymous):

so csc is 101/10?

OpenStudy (anonymous):

No, it's sqrt(101) over 10

OpenStudy (anonymous):

so (101)^2 over 10

OpenStudy (anonymous):

why is it squared

OpenStudy (anonymous):

 r^2 = x^2+y^2 And sqrt mean square root

OpenStudy (anonymous):

means*

OpenStudy (anonymous):

oo radical you mean, sorry about that that makesd sense now

OpenStudy (anonymous):

Just remember, when you do these kinds of problems, what x, y, and r stand for. sin = y/r cos = x/r tan = y/x csc = r/y sec = r/x cot = x/y r^2 = x^2+y^2